### v1

# Benchmarking Gaussian and non‑Gaussian input states with a hybrid sampling platform

Abstract
The original boson sampling paradigm—consisting of multiple single-photon input states, a large interferometer, and multi‑channel click detection—was originally proposed as a photonic route to quantum computational advantage. Its non‑Gaussian resources, essential for outperforming any classical system, are provided by single‑photon inputs and click detection. Yet the drive toward larger experiments has led to the replacement of experimentally demanding single‑photon sources with Gaussian states, thereby diminishing the available non‑Gaussianity—a critical quantum resource. As the community broadens its focus from the initial sampling task to possible real‑world applications, it becomes crucial to quantify the performance cost associated with reducing non‑Gaussian resources and to benchmark sampling platforms that employ different input states. To address this need, we introduce the Paderborn Quantum Sampler (PaQS), a hybrid platform capable of performing sampling experiments with eight Gaussian or non‑Gaussian input states in a 12‑mode interferometer within a single experimental run. This architecture enables direct, side‑by‑side benchmarking of distinct sampling regimes under otherwise identical conditions. By employing a semi-device-independent framework, offering certification that does not rely on prior knowledge of the interferometer or the input states, we verify that the observed data cannot be reproduced by any classical model—a prerequisite for demonstrating quantum advantage. Applying this framework, we observe clear performance gains arising from non‑Gaussian input states.
Boson sampling (BS), now often termed Fock‑state boson sampling, was first proposed as a route to demonstrating a provable quantum computational advantage [1, 2]. Early proof‑of‑concept experiments quickly followed, but their scale was limited by the difficulty involved in generating many high‑quality single‑photon inputs [3, 4, 5, 6]. Researchers began to explore the use of alternative input states—such as two-mode squeezed-vacuum (TMSV) states in the protocol now known as scattershot boson sampling (SBS)—as a means of reducing the experimental overhead [7]. Ultimately, the issue of system scaling was resolved when it was revealed that driving the system with single-mode squeezed-vacuum (SMSV) states preserves the computational complexity of the original proposal without the need for non-Gaussian input resources [8, 9]. The resulting protocol, Gaussian boson sampling (GBS), enabled a sequence of increasingly large experiments that pushed steadily toward a robust demonstration of quantum computational advantage [10, 11, 12, 13, 14, 15, 16].
Concurrently, researchers began investigating the utility of sampling architectures for addressing tasks with real‑world relevance. Notably, problems such as molecular vibronic spectroscopy [17] and a variety of graph‑theoretic computations were mapped onto GBS-type devices [18, 19, 20, 21, 22], and dedicated implementations successfully solved these tasks [14, 15]. It remained unclear, however, whether these problems could exhibit a genuine quantum computational advantage in this setting. Subsequent work revealed that many could, in fact, be solved efficiently using classical means [23, 24, 25, 26]. This realization has led to two key outcomes: it has motivated investigations into alternative applications such as Monte Carlo integration [27, 28, 29] and machine learning [30, 31, 32, 33], and it has stimulated deeper exploration of integrating additional non-Gaussian resources in these system to broaden their computational capabilities.
The non‑Gaussian element in GBS systems arises from photon‑number‑resolved (PNR) detection (or in earlier implementations, click detection). Extending the computational capabilities of such systems is therefore most easily achieved by supplementing this resource, either through advanced operations within the interferometer [31, 34, 35, 36], or by driving the interferometer with more exotic non‑Gaussian input states, whose impact has been the subject of extensive investigation [37, 38, 39, 40, 41, 42, 43, 44]. Consequently, the next generation of sampling devices will require a conceptual redesign of both their system architecture and resource allocation.
A corresponding rethink is necessary for benchmarking the performance of these newly conceived sampling systems. A commonly used verification technique is to quantify the closeness of generated data to that which is expected from theory, typically implemented using a metric known as the total variation distance [1]. However, this method does not serve as an ideal benchmarking tool for different sampling configurations, as it makes no direct statement about presence of quantumness in the system, particularly in the presence of system imperfections—a requirement to perform beyond classical limitations. Progressing the field therefore requires new certification tools capable of identifying genuinely non‑classical features—or quantumness—and experimental platforms that enable fair, side‑by‑side comparison of different architectures.
In this work, we introduce the Paderborn Quantum Sampler (PaQS), an experimental platform designed to benchmark multiple sampling schemes within a single experimental run. To compare the performance associated with different input‑state resources—both Gaussian and non-Gaussian—as illustrated in Figure 1, we further develop a benchmarking framework based on normally ordered moments of the photon‑number operator [46, 47, 48, 49, 50, 51]. In contrast to previous approaches, this method certifies the presence of quantumness in the measured data— a necessary prerequisite for any claim of quantum computational advantage [52].
Our results reveal stark differences between the generated GBS and SBS datasets. The quantumness of SBS data increases steadily with the mean photon number of the input states, whereas the GBS data displays strong non‑classical signatures at low mean photon numbers, but fails to maintain them as the brightness increases. This divergence underscores the fundamentally different behaviors arising from distinct input‑state resources and highlights the need for a conceptual shift when expanding sampling architectures to operate beyond their original task.
Figure 1: Conceptual implementations for various sampling configurations. In all cases, the input fields enter from the bottom of the setup and are detected with PNR detectors upon exiting the interferometer at the top. The input states of the interferometer can be either eight SMSV states, heralded Fock states, or thermal states, corresponding to implementing GBS, SBS, and thermal boson sampling (TBS), respectively. To implement SBS, the second mode of each TMSV input state is detected with PNR detectors to herald the number of photons entering each interferometer port. To realize sampling with thermal states - thermal boson sampling (TBS) - the second mode of each TMSV state is discarded, i.e., traced out.
Experimental Setup
Figure 2 presents a schematic of the PaQS system, divided into its constituent subsystems for clarity. The design of PaQS emphasizes integration: both the squeezed-light source (or parametric down-conversion source, PDC) and the programmable interferometer are realized on integrated photonic platforms. Notably, the integrated source, in conjunction with an electro-optic modulator and polarizing beamsplitter, enables dynamic switching between TMSV- and SMSV-state generation. This architecture allows multiple sampling configurations to be acquired during a single experimental run, thereby ensuring that results from the different sampling schemes are directly commensurable. Furthermore, the picosecond squeezed-light pulses generated by the waveguided source enables the implementation of intrinsic photon-number resolution (PNR) in the employed superconducting nanowire single-photon detectors (SNSPDs). The integrated interferometer offers convenient and complete programmability of the implemented unitary transformation, with an average insertion loss below 3 dB (see Supplementary Material). An overview of the key subsystems of PaQS is given in Figure 2, while detailed descriptions are provided in the Supplementary Material.
Figure 2: Schematic of the PaQS sampling system. The subsystems are separated to highlight the modular design. EOM — electro-optic modulator; PBS — polarizing beam splitter; SLM — spatial light modulator; GR — grating; PDC — parametric down-conversion source; F — filtering stage; $C$ — (temporal) compensation crystal; PLCU — path-length compensation unit.
A train of pulses is first selected from a recently developed Menlo laser system consisting of an ultra-low-noise stabilized frequency comb with a central wavelength of 1544 nm (Menlo Systems GmbH), a laser extension unit for providing control signals and laser amplification, and a frequency doubling stage to generate 772 nm light. This train of pulses is spectrally shaped to further optimize the pump for squeezed-light generation. The $\approx 2$ ps pulse duration squeezed-light pulses are produced in a single-pass, type-II periodically poled potassium titanyl phosphate (PPKTP) waveguide (AdvR Inc.), then spectrally filtered using a silicon window, to remove the pump filter, and a 2 nm bandpass filter, to remove the phase-matching side lobes. After filtering, the pulses are group-delay-compensated using a compensation crystal (CC) before passing through an electro-optic modulator (EOM2, QUBIG GmbH) and a polarizing beam splitter (PBS). By varying the voltage applied to EOM2, we can rapidly switch (rise/fall time of 4 ns) between generating two independent SMSV states and a TMSV state at the PBS output. This tunability enables the implementation of GBS [8], SBS [7], and TBS [53] configurations. One PBS output is directed into a multiplexed detection scheme comprising a splitting tree and eight SNSPDs (Single Quantum), while the other arm passes through an active time-to-space demultiplexer (QUBIG GmbH) and a path-length compensation system consisting of multiple path-length compensation units (PLCUs - Menlo Systems GmbH) that route the pulses into individual fibers and synchronize their arrival times.
The synchronized pulses subsequently interfere in a 12-mode, integrated, fully programmable interferometer (QuiX Quantum), of which eight channels are used. The photons present in each output channel are detected using eight SNSPDs (Single Quantum). Intrinsic photon-number detection up to three photons is realized on all detectors using a recently developed technique that determines photon number from the rise time of the SNSPD signal measured using a high resolution time tagger (Swabian Instruments GmbH) [54, 55]. Further experimental details are provided in the Supplementary Material.
System Characterization
We summarize below the principal methods used to characterize and verify the operation of the PaQS subsystems. The corresponding results are shown in Figure 3, with further details provided in the Supplementary Material.
The transmission of PaQS is quantified using the Klyshko method [56], applied individually to each signal mode. The system is configured for SBS operation in the low-gain regime, and an identity transformation is programmed on the interferometer. The Klyshko efficiency $\eta^{(i)}=C_{i}/H_{i}$ for each mode $i$ is given by the ratio of the measured coincidence rate $C_{i}$, between the relevant heralding bins and the output detector for that mode, and the single-detection rate $H_{i}$, in the heralding arm. The results, presented in Figure 3 a, show that all modes achieve efficiencies above $6.5\%$, with an average of $8.7\%\!\pm\!1.5\%$.
Single spectro-temporal-mode operation of our type-II waveguide source is verified through second-order autocorrelation measurements [57]. The system operates in SBS mode while photon statistics are recorded for one output of the PBS as the mean photon number is varied by adjusting the pump power driving the nonlinear process. As shown in Figure 3 b, the second-order correlation function converges to $g^{(2)}=1.95\pm 0.03$ at higher power, corresponding to an expected Schmidt mode number $K=1.05\pm 0.03$ effective modes—confirming both spectro-temporal purity and suitability for sampling experiments. The increase in $g^{(2)}$ at low mean photon numbers originates from a residual SMSV contribution because of imperfect polarization separation at the PBS (see Supplementary Material).
Next, we assess the ability of PaQS to switch between SMSV and TMSV generation. Operating in the low‑gain regime, we record the coincidence counts $n_{\mathrm{C}}$ between the herald and signal detectors as the driving voltage of EOM2’s high-voltage amplifier is varied (Figure 3 c). The normalized coincidence signal varies nearly sinusoidally with a fitted visibility of $V=96.3\%\!\pm\!1.2\%$ (See Supplementary Material), demonstrating precise and continuous tunability between SBS (maximum coincidences) and GBS (minimum coincidences) configurations. The difference in number of normalized coincidence counts for the two displayed SBS configurations (at 0 V and $\approx$1.35 V) is due to the different spatial-mode profiles of the two polarization modes within the waveguide.
Finally, Figure 3 d illustrates the precision with which we can tune the optical path-length difference between input modes using the PLCUs in order to optimize the indistinguishability between different sources (further details in the Supplementary Material). Implementing the SBS configuration in the low‑gain regime, we set the interferometer to realize a 50:50 beam splitter between any pair of chosen input modes. By varying the temperature of one PLCU fiber, we tune its optical length and record the resulting four‑fold coincidence counts $n_{\mathrm{Corr}}$ between the two beam-splitter outputs and their corresponding heralds, correcting for higher photon‑number contributions as described in the Supplementary Material. This configuration effectively realizes Hong–Ou–Mandel (HOM) interference between distinct sources [58, 59, 60]. Given the measured spectro-temporal mode number $K=1.05$ from Figure 3 b, the expected maximum visibility is $V\approx 95\%$. Some measurements approach this limit, while slight reductions in others are attributed primarily to imperfections in specific 50:50 beam splitters implemented in the interferometer.
[…]
Violation of Classicality Bounds
We now apply the developed benchmarking framework to experimental data produced by PaQS, which directly probes for the presence of nonclassical photon‑number correlations between output modes, or quantumness. This approach leverages the fact that quantumness is a prerequisite for any genuine quantum computation or for generating results that cannot be efficiently simulated [52]. The proposed bound is independent of the implemented unitary transformation, making the analysis semi-device-independent and therefore robust against certain experimental imperfections. Moreover, the ability of PaQS to switch rapidly between GBS and SBS/TBS enables a fair benchmarking between these sampling regimes.
Our nonclassicality criterion is grounded in Glauber’s coherence theory [61]. Recasting matrix of moments approaches [47, 48], we identify quantumness in the photon‑number covariances $\mathrm{Cov}(n_{j},n_{k})=\langle\hat{n}_{j}\hat{n}_{k}\rangle-\langle\hat{n}_{j}\rangle\langle\hat{n}_{k}\rangle$ between output modes $j,k$. For classical light, the MM‑mode matrix $(\mathrm{Cov}(n_{j},n_{k})-\delta_{j,k}\langle\hat{n}_{j}\rangle)_{j,k=1,\ldots,M}$, where $\delta_{j,k}$ denotes the Kronecker delta, is positive semidefinite. Therefore, observation of a negative eigenvalue constitutes unambiguous evidence of nonclassical photon‑number correlations and, therefore, the presence of quantumness in the system.
We apply this measure to our experimental datasets, as summarized in Figure 4. Panel a shows the minimum eigenvalues obtained for TBS, GBS, and SBS data generated in a single measurement run with a mean photon number of $0.569\pm 0.011$, together with GBS data taken at a higher mean photon number of $2.152\pm 0.090$. During each $\sim$11‑minute run, the system alternates between GBS and SBS/TBS configurations every 20 s to mitigate long‑term drifts. Minimum eigenvalues are computed from 1 s data blocks, each encompassing one million samples. This temporal segmentation enables investigation of possible time‑dependent effects, which may arise due to drifts in the phases of the input states. As expected, we find that only the GBS data exhibits such a temporal dependence due to the phase‑space asymmetry of SMSV states - i.e. the recorded minimum eigenvalues vary between different measurement bins as the input phases evolve. Notably, the GBS data at $\langle n\rangle=0.569$ shows that more than 80% of the recorded minimum eigenvalues are negative — thereby verifying quantumness in the generated data, whereas at $\langle n\rangle=2.152$, the system displays quantum correlations less than 2% of the time.
Figure 4: Quantumness analysis. a, Recorded minimum eigenvalues during a full measurement run at a mean photon number of 0.569, together with GBS data at 2.152 for comparison. Each data point corresponds to a 1 s integration window; the grey band denotes the uncertainty. b, Highest and lowest observed minimum eigenvalues for generated GBS, SBS, and TBS data across all measured mean photon numbers. Shaded regions represent uncertainties of reported values and have been connected to guide the eye. Uncertainties are estimated from counting errors.
To benchmark the effect of increasing the input energy between the various sampling regimes, Figure 4b plots the obtained minimum eigenvalues for GBS, SBS, and TBS as a function of mean photon number. The range of these values at each brightness is illustrated by plotting both the lowest and highest obtained minimum eigenvalues. For GBS at $\langle n\rangle=2.152$, these extrema correspond to the highest and lowest eigenvalues from the data in Figure 4a. Both GBS and SBS data exhibit clear quantum signatures, whereas TBS data, as expected, shows no violations of the classical bound.
Discussion
We begin by examining the dependence of the input state brightness on the measured degree of quantumness. For TBS (Figure 4 dark and light blue lines), the minimum eigenvalue increases with mean photon number and, at the highest brightness, rises above the nonclassicality bound by more than $13\,\sigma$. This behavior is expected: increasing brightness simply adds photons from a thermal distribution, increasing photon‑number variance. In contrast, the SBS data (Figure 4 dark and light green lines) contains minimum eigenvalues that continuously decrease with increasing input brightness. At the highest mean photon number recorded, SBS data violates the classicality bound by more than $36\,\sigma$, therefore, brighter input states are advantageous in the SBS configuration.
In the GBS configuration, however, increasing the brightness of the input SMSV states does not yield a stronger or more statistically significant violation of the classical bound beyond a certain limit (Figure 4 dark orange line). The maximum violation exceeds $26\,\sigma$ at an average photon number of $\langle n\rangle=0.172$, while the local minimum at $\langle n\rangle=0.569$ corresponds to a $14\,\sigma$ violation. Once the mean photon number exceeds the value at this local minimum, the system’s quantumness no longer continues to increase. Furthermore, Figure 4 reveals that at a mean photon number of $\langle n\rangle=2.152$, the two-mode correlators from the GBS data do not exhibit quantumness for over 90% of the input phase values. Therefore, a classical model could efficiently reproduce this data.
Understanding the cause for the observed reduction in quantumness is paramount. One natural hypothesis is that the reduced quantumness results from increased input-state impurity caused by system losses, consistent with recent work that decomposes impure SMSV states into a classical thermal component and a pure squeezed component [25]. Simulations performed using The Walrus [62] library, however, reveal that this behavior is compatible with simulations based on a lossless system (see Supplementary Material), asking for explanations beyond input state impurity. Furthermore, the simulations were run with fixed input phases for each run and were calculated up to a photon number of 15—also ruling out the possibility that the reduction in quantumness stems from phase drifts or limited PNR.
We therefore postulate an alternative source for the observed behavior—that as the average photon number of the input states increases, the entanglement present at the output of the system begins to reside in higher-order correlations—the analysis of which would require experimental PNR resolution of greater than 4 and would require a significant experimental overhead. This effect is likely not seen in the SBS data because heralding collapses the photon-number superposition present in GBS and instead defines a well known input photon-number distribution. A deeper investigation of this effect is outside the scope of this paper and is left for future work. Although the precise mechanism remains unresolved, our findings demonstrate firstly that SBS outperforms GBS in our experimental implementation and furthermore, that there likely exists an optimal level of squeezing in GBS systems—an observation previously discussed in related contexts [14] but deserving renewed attention in light of the present methodology. These results demonstrate that modifying the input state can yield markedly different system behavior and introduces a new framework for benchmarking emerging platforms aimed at extending the computational capabilities of sampling systems.
Methods
Pump Source
Pump pulses are generated by a newly developed Menlo laser system that begins with a SmartComb system producing femtosecond pulses in the telecom C‑band. Light from the SmartComb enters a laser extension unit that provides optical gain and generates various control signals. Finally, the amplified pulses are frequency‑doubled in a free‑space second‑harmonic stage to yield picosecond pulses centered around 772 nm with a full-width at half maximum bandwidth of approximately 1 nm, a repetition rate of 80 MHz and pulse energies up to approximately 12 nJ.
[…]
Squeezed‑State Generation
The prepared pump‑pulse train enters the squeezed‑light source, which comprises a waveguided parametric down‑conversion (PDC) source, a temporal compensation crystal (CC), EOM2, a PBS, and fiber‑coupling optics. The 20 mm‑long waveguided PPKTP crystal (AdvR Inc.) is designed for a type‑II PDC process that creates signal and idler fields that are spectrally indistinguishable and that generates ps-long squeezed light pulses in a single spectro-temporal mode.
Following the waveguide, the output passes through a filtering stage composed of a silicon window that suppresses the pump and a 2 nm spectral filter that removes phase matching side lobes. An 8.8 mm‑long KTP group‑delay compensation crystal (CC) then corrects for birefringent walk‑off accumulated in the waveguide, ensuring temporal overlap of the signal and idler pulses. The filtered, temporally synchronized fields subsequently enter an electro‑optic modulator (EOM2, QUBIG GmbH HVOS‑SWIR‑1.5k100), which sets the output polarization of the transmitted modes on a polarizing beam splitter (PBS). The PBS output ports are fiber‑coupled. Interference of the two polarized fields at the PBS produces independent SMSV states used for GBS implementations, whereas deterministic separation of the fields yields a TMSV state used to realize SBS or TBS configurations in a single experimental data run.
Multiplexed Detector and Heralding
A multiplexed detection module attached to one PBS output both spatially and temporally separates the incoming pulses to measure the photon number in this heralding arm. This module enables photon‑number heralding for implementing higher-order SBS and allows verification of the lack of correlations between the two PBS outputs in the GBS configuration. Discarding the heralding component of the data obtained SBS effectively prepares thermal input states for the interferometer [57].
The multiplexed architecture is required because, in the absence of a second time‑to‑space demultiplexer, the 12 ns pulse separation is shorter than the SNSPD dead time of approximately 80 ns. The implemented scheme minimizes the probability of photon arrival during detector dead time, achieving an average heralding efficiency of $38.4\,\%\!\pm\!0.4\,\%$.
Time‑to‑Space Demultiplexing
Light exiting the second PBS output passes through an active time‑to‑space demultiplexer (DMX16_80, QUBIG GmbH), which routes each input pulse into individual output fibers via an EOM switching tree with an efficiency exceeding 80%. Each output fiber is fed into a temperature‑stabilized path‑length compensation unit (PLCU, Menlo Systems GmbH) consisting of segments of fiber of variable length. These include approximately 60 m of fiber wound around a piezo actuator and additional lengths compensating the 12.5 ns time difference between consecutive pulses. The optical path length can be tuned using temperature control and mechanical strain.
The phase of the fields exiting the PLCUs is currently free‑running; no active phase stabilization is implemented. Nonetheless, each PLCU can serve as an actuator for future active control once an appropriate error signal is defined. For the present implementation, all phase‑sensitive components are enclosed in polyethylene foam housing to achieve high passive stability.
Integrated Interferometer
The temporally synchronized pulses emerging from the demultiplexer are fed via separate fibers into a 12‑mode integrated interferometer (QUIX Quantum) that provides full programmability of the implemented unitary transformation. This interferometer is realized in the low-loss silicon nitride platform and exhibits an average insertion loss of $2.87\!\pm\!0.37$ dB. The mean similarity between implemented and target unitary matrices drawn from a Haar‑random ensemble is approximately 94% for 12‑dimensional and 97% for 8‑dimensional transformations.
Photon‑Number Detection
Photon numbers in the eight utilized output modes of the interferometer are measured using eight SNSPDs (Single Quantum) with quantum efficiency exceeding 90% and timing jitter below 20 ps. An intrinsic PNR scheme is implemented, relying on the $\approx 2$ ps pulse duration of the squeezed‑light states and the sub‑2 ps jitter of the Swabian Instruments Time Tagger X. Using this technique, our implementation discriminates events with up to three photons per pulse with high confidence [55, 63].
Data Availability
Source data are provided within the paper. The experimental data used in this paper will be made publicly available in a Zenodo repository on publication.
Acknowledgments
M.S would like to acknowledge many fruitful discussions with various colleagues and friends: Sevag Gharibian, Dhruva Sambrani, Hamid Naeij, Tina Chen, Dorian Rudolph, Jonas Kamminga, Robert Schade, Nathan Walk, Helen Chrzanowski, Jens Eisert, Dario Cilluffo, Martin Plenio, Andrew White, Fatemeh Mohit, Craig Hamilton, Shivani Singh, Aurèl Gàbris, Magdalena Paryzkova, Igor Jex, Tim Ralph, Ryan Marshman, Nibedita Swain.
F.L would like to thank Laura Serino for help with the SLM.
F.S. is part of the Max Planck School of Photonics supported by the Dieter Schwarz Foundation, the German Federal Ministry of Research, Technology and Space (BMFTR), and the Max Planck Society.
T.S and T.B acknowledge partial funding by the European Union (ERC, QuESADILLA, 101042399). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
This work has received funding from the German Federal Ministry of Research, Technology and Space within the PhoQuant project (Grant No. 13N16103).
Author contributions
M.S. and K.-H.L. designed the experiment with input from B.B, and C.S.; K.-H.L. and S.A. built the experimental setup with the help of F.L., J-L.E., C.P., F.Sc., J.L., M.S., M.Y., A.K., C.S., M.L., O.K., and F.St.; M.Y., A.K., and R.H. developed and/or partially installed the laser system and path length compensation units; F.Sc. designed and constructed the multiplexed detection scheme for heralding with aid from J.L.; M.Y., S.A., K.-H.L. implemented the PLCU scheme; F.Sc., T.S., T.B., J.L. developed and implemented the PNR detection scheme; F.L. implemented the pulse shaper with help from K.-H.L., and F.Sc.; K.-H.L. designed and built the squeezed-light source with the help of F.L., J-L.E., and F.Sc.; J.-L.E., K.-H.L., S.A, M.R. performed the experimental runs and data collection with the help of J.L. and C.P.; J.-L.E. completed data analysis with the help of L.A., J.S., J.L., and C.P.; M.S. prepared the manuscript with the help of K.-H.L., S.A., J-L.E., F.L., C.P., F.Sc., L.A., J.S., M.R., B.B., C.S., and J.L.; All authors discussed the results and reviewed the manuscript.
Competing interests
The authors declare no competing interests.
Additional information
Supplementary information
Correspondence and requests for materials
should be addressed to Michael Stefszky, Jan-Lucas Eickmann, or Christine Silberhorn.
Data and material availability:
All data are available in the manuscript, online at address or in the supplementary material.